3. Two companies Toxic Wastes Management and Sludge Industries both pollute a nearby lake Each firm dumps 1,000 gallons of goo every day into the lake. As a consequence the lake has lost its clarity and the fish are dying. Local residents want to see the lake restored. Toxic Wastes production process depends heavily on being able to dump the goo into the lake. It would cost the Toxic Waste $10 per gallon to clean up the goo it creates. Sludge Industries can clean up its goo at a cost of $2 per gallon. a. If the local government cuts the legal goo emissions in half for each firm what are the costs to each firm to comply with the law? What is the total cost of both firms in meeting the goo emissions standard? b. Another way of cutting goo emissions in half is to assign each firm tradable pollution permits that allow 500 gallons of goo to be dumped into the lake every day. Will each firm still dump 500 gallons of goo? What are the costs to each firm under tradable pollution permits? What is the total cost of both firms under the pollution permit plan?
Answer #1

Answer 3 (a) : Emission standards and toxic emission control play a critical role in the upkeep of the economy of a country both directly and indirectly. If the pollution hazards are not dealt with implicitly, it can in due coarse of time lead to the total extinction of the social beings and natural factors of production thereby directly impact the economic downfall of a nation. In this case, both Toxic Wastes Management and Sludge industries are directly impacting to the pollution increase in the locality they reside by dumping the goo in to the lake. If the local Government cuts the legal goo emission in half for each firm, both the firms would be directly impacted by the new law. As per the given condition, Toxic Wastes Management production is highly dependent on the ability to dump the loo in the lake and it would cost the Toxic Wastes management $10 per gallon to clean up the goo it creates, therefore , if the law enforces to cut the goo emission to half , then it would also cost Toxic Wastes management half the amount of total money that it would spend in the clean up of the goo, i.e.

Total cost of clean up of the goo by Toxic Wastes Management per day = 1000 gallons x $ 10 (cost of cleanup of one gallon)

                                                                                                                                  = $ 10000 per day

Now, if the emission law mandates to cut the waste dumping to half, then it would cost half this total cost per day, i.e. = 10000 / 2 = $ 5000 per day.

Similarly, Total cost of cleanup of the goo by Sludge Industries per day = 1000 gallons x $ 2 (cost of cleanup of one gallon)

                                                                                                                                  = $ 2000 per day

Now, if the emission law mandates to cut the waste dumping to half, then it would cost half this total cost per day, i.e. = 2000 / 2 = $ 1000 per day.

                                If both firms unite in meeting the new goo emission standards, then the total cost of clean up of the goo would be = $ 5000 + $1000 divided by 2, i.e. = 6000/ 2

                                                                                                                    = $ 3000 per day for both firms combined.

Answer 3 (b) : If the tradable pollution permit of allowing 500 gallons of goo per day per company is implemented, then it would comparatively be more advantageous for both the firms. Since both the firms are now allowed 500 gallons per day and the Sludge Industries not over dependent in the dumping of the goo in the lake , most likely due to the less production by the industry would not require the complete quota of 500 gallons and the Toi castes management may come in to a contract with the Sludge industries to buy of the remaining of its quota and use it for itself. It will reduce the cost of bot the firms , but it will ultimately not reduce the waste disposal in to the lake by much.

Total cost of Tradable pollution permit by Toxic Wastes Management per day = 500 gallons x $ 10 (cost of cleanup of one gallon)

                                                                                                                                  = $ 5000 per day

It will ask the Sludge industries to buy the remaining 300 gallons of quota of the disposal at a reduced rate= 5000 + 1000 , i.e.total cost would be $6000/2 = $ 3000 per day

Similarly , Total cost of Tradable pollution permit by Sludge industries per day = 200 gallons x $ 2 (cost of cleanup of one gallon)

                                                                                                                                  = $ 400 per day

                                If both firms unite in meeting the new goo emission standards, then the total cost of clean up of the goo would be = $ 3000 + $600 divided by 2, i.e. = 3600/ 2

                                                                                                                    = $ 1800 per day for both firms combined.